A manufacturer of galvanized chains advertises the load capacity of a 16-foot length of chain has an average of 2650 lbs. and a standard deviation of 520 lbs. If a retailer receives an order of 75 chains, what is the probability that the sample mean will be greater than 2,700 lbs?
Answer by stanbon(75887)

A manufacturer of galvanized chains advertises the load capacity of a 16-foot length of chain has an average of 2650 lbs. and a standard deviation of 520 lbs. If a retailer receives an order of 75 chains, what is the probability that the sample mean will be greater than 2,700 lbs?
---
t(2700) = (2700-2650)/[520/sqrt(75)] = 0.8327
----
P(x-bar > 2700) = P(t > 0.8327 when df = 74) = 0.2038
==========================
Cheers,
stan H.